OpenStudy (anonymous):

A circle with a diameter of 8 inches has a circumference of?

OpenStudy (anonymous):

There's a formula for the Circumference of a Circle. Its 2 times Pi (3.14..) times the Radius

OpenStudy (anonymous):

The Radius is equals to the diameter divided by 2.

OpenStudy (anonymous):

so 2*3.14*4 = circumference?

OpenStudy (anonymous):

Correct. Exactly that. Now if they ask you for an exact answer, leave Pi as Pi, then just multiply 2*4

OpenStudy (anonymous):

it says answer as a multiple of pi or use 3.14

OpenStudy (anonymous):

oh, well in that case you can use either. It would be \[8\pi \] or 25.12

OpenStudy (anonymous):

ok thanks! how would i find the area?

OpenStudy (anonymous):

There's a formula for that as well. It is (Pi) * r^2 (or r*r) where r is the radius. Or diameter divided by 2

OpenStudy (anonymous):

\[\pi r^2\] to be exact

OpenStudy (anonymous):

so 50.24 in.

OpenStudy (anonymous):

yep, correct

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no problem